Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

probability

Expert replies
by CITI29 » Wed Jul 23, 2008 8:20 pm
If a six sided die is rolled three times, what is the probability of getting at least one even number and at least one odd number?
1/8 1/4 1/2 3/4 7/8
Join the discussion
Source: — Problem Solving |

Re: probability

by Stuart@KaplanGMAT » Wed Jul 23, 2008 8:39 pm
CITI29 wrote:If a six sided die is rolled three times, what is the probability of getting at least one even number and at least one odd number?
1/8 1/4 1/2 3/4 7/8
This question is much more complicated than it needs to be, so let's simplify it.

50% of the time it will be even, 50% of the time it will be odd. So, we can pretend that the die has two sides (or even pretend that this is a coin flip question).

We could do this by brute force. If we roll 3 times, the diff events are:

EEE
EEO
EOE
EOO
OOO
OOE
OEO
OEE

6 of those events have at least one E and at least one O, so the answer is 6/8 = 3/4.

We could also do this using probability.

Prob = # desired events/total # of possibilities

Also, we know that Prob (what you want) = 1 - prob (what you don't want)

The "1 - what you don't want" approach is often the quickest way to attack tough probability questions on the GMAT.

Since this is a binary scenario (i.e. each option happens 50% of the time), the total # of possibilities is going to be 2^n, where n is the number of events. In this case, that's 3.

So, there are 2^3 = 8 total possibilities.

The only 2 events that we do NOT want are OOO and EEE. So, we don't want 2/8 = 1/4 of the things to happen.

So, Prob(do want) = 1 - 1/4 = 3/4

The second solution (using the formula) is more elegant, but brute force is probably quicker. Remember, on test day you don't get any bonus points for the beauty of your solution, we want fast and furious!
Image

Stuart Kovinsky | Kaplan GMAT Faculty | Toronto

Kaplan Exclusive: The Official Test Day Experience | Ready to Take a Free Practice Test? | Kaplan/Beat the GMAT Member Discount
BTG100 for $100 off a full course
Join the discussion